OpenStudy (anonymous):

Change the following equation into slope intercept form. –3x + 2y = 8

OpenStudy (anonymous):

\[2y = 8 + 3x\] \[y = \frac{ 3 }{ 2 }x + 4\]

OpenStudy (anonymous):

Thank you so much :)